OpenStudy (anonymous):

400 raffle tickets are to be sold for $2.00 each. there will be one prize of $200 awarded, and 2 prizes of $50.00. calculate the expected value for this game. If the game is not fair calculate a ticket price to make it fair.

OpenStudy (anonymous):

400 * 2 = 800 - 200 = 600 - 100 = 500 $500 is the expected value To make it fair the ticket price should be $300 / 400 = $0.75